A tractor manufacturing company knowa that the number of heavy duty tractors it sell depends on price per unit. If the demand function is
P=25000-2q and cost of production is
Tc=2000000-20000q +5q2
Where q :number of units sold
P:price
Tc : total cost
How many units should be produced to maximize profit and what is the maximum profit
1
Expert's answer
2021-09-29T17:14:20-0400

"P_d = 25000-2q\\
T_c =2000000-20000q +5q²"


marginal cost = 10q -20000

marginal revenue = -2


In order to maximize profit, the firm should produce where its marginal revenue and marginal cost are equal.


"10q -20000 = -2\\

q = 2000"